OpenStudy (anonymous):

Margaret had 5 3/8 yards of fabric to make curtains. She only used 4 3/4 yards. How much fabric was left over?

OpenStudy (anonymous):

4 3/4=4 6/8 = 38/8 5 3/8 = 43/8 43/8 - 38/8 = 5/8